# Quantum-Mechanical Suppression of Gas Accretion by Primordial Black Holes

### v1

@@ Abstract
The Schwarzschild radii of primordial black holes (PBHs) in the mass range of $6\times 10^{14}~{}{\rm g}$ to $4\times 10^{19}~{}{\rm g}$ match the sizes of nuclei to atoms. I discuss the resulting quantum-mechanical suppression in the accretion of matter by PBHs within gaseous astrophysical environments.
@@ 2 Quantum Suppression of Astrophysical Accretion
The gravitational and electric forces are of equal magnitude for a PBH mass of $M=3.8\times 10^{15}~{}{\rm g}$, and gravity wins at larger masses. The atomic radius in equation (3) shrinks below the Schwarzschild radius in equation (1) for a PBH mass $M=3.7\times 10^{17}~{}{\rm g}$. At this mass, the gravitational binding energy in equation (4) approaches a quarter of the electron rest-mass energy. For higher PBH masses, the Compton wavelength of the electron, $(\hbar/m_{e}c)=4\times 10^{-11}~{}{\rm cm}$, is smaller than the Schwarzschild radius, and the ground state of the bound electron can be inside the horizon.
For lower mass PBHs, accretion could be significantly suppressed in common astrophysical environments, where gas particles have non-relativistic thermal speeds and the mass density is much lower than the nuclear density of neutron stars. In the rarefied environments of the interstellar or intergalactic media, a low-mass PBH would accrete one particle at a time with a significant quantum-mechnical suppression in accretion rate relative to hydrodynamic expectations.
3 Discussion
@@
A bound electron can be captured by an atomic nucleus even without a central black hole. This well-known process, called electron capture, involves the absorption an electron from an inner atomic shell by a proton-rich nucleus of a neutral atom (Alvarez, 1937). The rate for this capture is related to the overlap of the electron wave function with the volume of the atomic nucleus.
@@
Including Hawking radiation (Hawking, 1974) would further lower the accretion rate since the Hawking temperature, $T_{\rm H}=(\hbar c^{3}/8\pi k_{B}GM)=2.7\times 10^{6}~{}{\rm eV}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-1}$, exceeds the binding energy $1{\rm{\tilde{R}y}}$ in equation (4), and the Hawking luminosity, $L_{\rm H}=(\hbar c^{6}/15360\pi G^{2}M^{2})=1.5\times 10^{26}~{}{\rm eV~{}s^{-1}}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-2}$, exceeds $1{\rm{\tilde{R}y}}/\tau_{1/2}$. The outward flux of high energy photons and electron-positron pairs could suppress accretion altogether in rarefied astrophysical environments.
@@
At high enough plasma densities, a bound state could involve multiple electrons and protons simultaneously. When the inflow of fresh protons into the PBH exceeds $\tau_{1/2}^{-1}$, the PBH could be charged positively up to a maximum charge, $Q_{\rm max}\sim(GMm_{p}/e^{2})e=(M/2.1\times 10^{12}~{}{\rm g})e$, at which the electric repulsion equals the gravitational attraction for external protons. At this maximum charge (which will be screened over the Debye length in the surrounding plasma), the accretion rate can be enhanced relative to equation (7) since the effective Bohr radius for the ground state of the innermost electron at $Q_{\rm max}$ shrinks below the Schwarzschild radius for PBH masses $M\gtrsim 8.6\times 10^{15}~{}{\rm g}$.
@@ Acknowledgements
This work was supported in part by Harvard’s Black Hole Initiative, which is funded by grants from JFT and GBMF.

### v4

# Quantum-Mechanical Suppression of Accretion by Primordial Black Holes

@@ Abstract
The Schwarzschild radii of primordial black holes (PBHs) in the mass range of $6\times 10^{14}~{}{\rm g}$ to $4\times 10^{19}~{}{\rm g}$ match the sizes of nuclei to atoms. I discuss the resulting quantum-mechanical suppression in the accretion of matter by PBHs in dense astrophysical environments, such as planets or stars.
@@ 2 Quantum Suppression of Astrophysical Accretion
The gravitational and electric forces are of equal magnitude for a PBH mass of $M=3.8\times 10^{15}~{}{\rm g}$, and gravity wins at larger masses. The atomic radius in equation (3) shrinks below the Schwarzschild radius in equation (1) for a PBH mass $M=3.7\times 10^{17}~{}{\rm g}$. At this mass, the gravitational binding energy in equation (4) approaches a quarter of the electron rest-mass energy. For higher PBH masses, the Compton wavelength of the electron, $(\hbar/m_{e}c)=4\times 10^{-11}~{}{\rm cm}$, is smaller than the Schwarzschild radius, and the ground state of the bound electron can be inside the horizon. The criterion that accretion of an electron is suppressed is identical, up to a numerical factor of order unity, to the condition that electron-positron pairs are emitted through Hawking evaporation of the black hole.
For lower mass PBHs, accretion could be significantly suppressed in common astrophysical environments, where gas particles have non-relativistic thermal speeds and the mass density is much lower than the nuclear density of neutron stars (Giffin et al., 2022). In the rarefied environments of the interstellar or intergalactic media, a low-mass PBH would accrete one particle at a time with a significant quantum-mechanical suppression in accretion rate relative to hydrodynamic expectations.
3 Accretion Rate
@@
3.1 Rarefied Cold Gas with Bound Electrons
The accretion of particles in bound states around a PBH or in tightly packed systems is different from absorption of free particles by a PBH (Unruh, 1976). A bound electron can be captured by an atomic nucleus even without a central black hole. This well-known process, called electron capture, involves the absorption an electron from an inner atomic shell by a proton-rich nucleus of a neutral atom (Alvarez, 1937). The rate for this capture is related to the overlap of the electron wave function with the volume of the atomic nucleus.
@@ 3.1 Rarefied Cold Gas with Bound Electrons
At high enough infall rate, a bound state could involve multiple electrons and protons simultaneously. When the inflow of fresh protons into the PBH exceeds $\tau_{1/2}^{-1}$, the PBH could be charged positively up to a maximum charge, $Q_{\rm max}\sim(k_{\rm B}T/m_{p})(GMm_{p}/e^{2})e=(T/1.1\times 10^{13}~{}{\rm K})(M/2.1\times 10^{12}~{}{\rm g})e$, at which the electric potential near the horizon equals the thermal energy for external protons at the background temperature $T$ (Medvedev & Loeb, 2024). At this maximum charge (which will be screened over the Debye length in the surrounding plasma), the accretion rate can be enhanced relative to equation (7) since the effective Bohr radius for the ground state of the innermost electron at $Q_{\rm max}$ shrinks below the Schwarzschild radius for PBH masses $M\gtrsim 8.6\times 10^{15}~{}{\rm g}$. However, at the high plasma temperatures expected near the PBH horizon, the electrons will not stay in bound states and the accretion rate will be dictated by the dynamics of free particles.
3.2 Dense Hot Plasma with Free Electrons
For a thermal plasma of free (unbound) particles, the accretion rate can exceed the value expected from equation (7) if the interparticle distance $d=(4\pi n_{e}/3)^{-1/3}$ is shorter than ${\tilde{r}}_{\rm B}$, where $n_{e}$ is the electron number density in the plasma. In this regime, the fastest accretion will involve one electron-proton pair per $\tau_{1/2}$, where the length scale ${\tilde{r}}_{\rm B}$ in equation (5) is replaced by the interparticle distance $d$ and the velocity scale ${\tilde{v}}_{\rm B}$ is replaced by the thermal speed of the particles $v_{\infty}$ far from the PBH (assuming a stationary PBH relative to the plasma rest frame). This substitution gives a net accretion rate of,
${\dot{M}}={4\over 3}\left({r_{\rm Sch}\over d}\right)\rho_{\infty}v_{\infty}\left(\pi r_{\rm Sch}^{2}\right)~{},$ (8)
where $\rho_{\infty}=m_{p}n_{e}$ is the plasma mass density far from the PBH.
It is instructive to compare this suppressed accretion rate to the Bondi accretion rate (Bondi, 1952), derived in the fluid approximation,
${\dot{M}}_{\rm Bondi}\approx\left({c\over v_{\infty}}\right)^{2}\rho_{\infty}v_{\infty}\left(\pi r_{\rm Sch}^{2}\right)~{}.$ (9)
The resulting suppression factor for $d>r_{\rm Sch}$ is,
$F_{\rm supp}=\left({{\dot{M}}\over{\dot{M}}_{\rm Bondi}}\right)\approx\left({v_{\infty}\over c}\right)^{2}\left({4\over 3}{r_{\rm Sch}\over d}\right)~{}.$ (10)
For $r_{\rm Sch}\ll d$ and $v_{\infty}\ll c$, we find that $F_{\rm supp}\ll 1$.
This result applies to high matter densities at which atoms are densely packed and $d<\tilde{r}_{\rm B}$. Such densities are relevant when PBHs pass through the interiors of planets or stars (Hawking, 1971; Luo et al., 2012; Caplan et al., 2024), with the most dense medium realized for PBHs trapped in the interior of a neutron star (Pani & Loeb, 2014; Giffin et al., 2022). For context, the Bondi accretion rate of a trapped PBH at the mean density of a white dwarf is substantial for $v_{\infty}<5\times 10^{3}~{}{\rm km~{}s^{-1}}$,
${\dot{M}}_{\rm Bondi}\approx 1.5\times 10^{5}~{}{\rm g~{}s^{-1}}\left({\rho_{\infty}\over 10^{6}~{}{\rm g~{}cm^{-3}}}\right)\left({M\over 3.6\times 10^{19}~{}{\rm g}}\right)^{2}\left({v_{\infty}\over 5\times 10^{3}~{}{\rm km~{}s^{-1}}}\right)^{-3}.$ (11)
Expressing this equation as ${\dot{M}}=AM^{2}$ with $A$ being a constant, yields a solution for the final PBH mass $M_{f}$ in terms of its initial mass $M_{i}$ as, $M_{f}=1/(1/M_{i}-At)$. This solution implies that over a timescale shorter than 10 Myr, the mass of a trapped PBH with $M_{i}=3.6\times 10^{19}~{}{\rm g}$, would diverge through Bondi accretion well above its initial value. Consequently, a trapped PBH could have had dramatic observable consequences on the energy budget or the ultimate fate of a white dwarf. However, in reality the PBH growth would be suppressed by the quantum-mechnical effects discussed above as well as by the Eddington limit on the accretion rate, which is $\sim 3\times 10^{5}~{}{\rm g~{}s^{-1}}(\epsilon/0.01)^{-1}(M/3.6\times 10^{19}~{}{\rm g})$, where $\epsilon$ is the radiative efficiency.
Other dense environments involve the thermal electron-positron plasma of the early universe (Curd et al., 2024), where the thermal speed of the particles is $\sim c$. The de Broglie wavelength of relativistic electrons and positrons in this thermal plasma of free particles is of order $d$, smaller than their Compton wavelength. This is properly incorporated in the suppression factor of equation (10) which is effective for $d>r_{\rm Sch}$. For a wavefunction with a de Broglie wavelength larger than the Schwarzschild radius, the relevant momentum is calculated far from the horizon, where most of the probability distribution of finding the particle resides.
4 Concluding Remarks
Including Hawking radiation (Hawking, 1974) would further lower the accretion rate since the Hawking temperature, $T_{\rm H}=(\hbar c^{3}/8\pi k_{B}GM)=2.7\times 10^{6}~{}{\rm eV}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-1}$, exceeds the binding energy $1{\rm{\tilde{R}y}}$ in equation (4), and the Hawking luminosity, $L_{\rm H}=(\hbar c^{6}/15360\pi G^{2}M^{2})=1.5\times 10^{26}~{}{\rm eV~{}s^{-1}}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-2}$, exceeds $1{\rm{\tilde{R}y}}/\tau_{1/2}$. The outward flux of high energy photons and electron-positron pairs could suppress accretion altogether.
The accretion of matter into asteroid-mass PBHs is only significant in the dense interiors of planets or stars, but can be ignored at the much lower cosmic-mean-density of matter and radiation in the Universe (Rice & Zhang, 2017).
@@ Acknowledgements
I am grateful for insightful comments from an anonymous referee. This work was supported in part by Harvard’s Black Hole Initiative, which is funded by grants from JFT and GBMF.